11. A sample of 62 aircraft “C” checks generated an average of 462 write-ups (squawks), with a standard deviation of 57. The VP of operations wants to know how many squawks to expect on the next “C” check performed. Can you give him an estimate based on a 90% degree of confidence?
Answer #1

solution:

Given that,

Point estimate = sample mean = \bar x = 462

Population standard deviation =   \sigma = 57

Sample size = n =62

At 95% confidence level the z is ,

\alpha = 1 - 95% = 1 - 0.95 = 0.1

\alpha / 2 = 0.1 / 2 = 0.05

Z\alpha/2 = Z0.05 = 1.645


Margin of error = E = Z\alpha/2 * ( \sigma /\sqrtn)

= 1.645* (57 /  \sqrt62 )

= 11.9081
At 90% confidence interval estimate of the population mean
is,

\bar x - E < \mu < \bar x + E

462 -11.9081 <  \mu < 462+ 11.9081

450.0918 <  \mu < 473.9081

(450.0918 , 473.9081)